a train a mile and a half long takes 1 minute and a half to go through a tunnel a mile and a half long. how fast is it going?

Answer by stanbon(57962)

a train a mile and a half long takes 1 minute and a half to go through a tunnel a mile and a half long. how fast is it going?
--------------------------------
distance the back of the train goes from the time the front enters
the tunnel to the time that the back exits the tunnel:
= 1.5 + 1.5 miles
--
time to do that = 1 min = (1/60) hr.
-----
Equation:
rate = distance/time
---
rate = 3 miles/(1/60) hr = 180 mph
============================
Note: If you just want the distance the front goes from the
time it enters the tunnel till the front exits the tunnel
the distance is 1.5 miles
---
rate = 1.5(1/60) = 90 mph
============================
Cheers,
Stan H.